write each dec__al as a common fraction in lowest terms
A)0.075 b)0.75

Answer by stanbon(75887)

write each dec__al as a common fraction in lowest terms
A)0.075
Ans: 75/1000 = 3/40
---------------------
b)0.75
Ans: 75/100 = 3/4
=========================
Cheers,
Stan H.